OpenStudy (anonymous):

Write the complex number in polar form with argument θ between 0 and 2π. − 5πi

OpenStudy (dumbcow):

\[z = 0-5\pi i\] plot the point  \[\rightarrow 5\pi(\cos \frac{3\pi}{2} -i \sin \frac{3\pi}{2})\]